OpenStudy (anonymous):

two masses 9 kg ang 4kg are placed at a distance of 10m from each then the value og potential where gravitational field is zero??

OpenStudy (anonymous):

i m not gettin formula for this .

OpenStudy (stormfire1):

The only formula you need is \[Fg = G \frac{m _{1}m _{2}}{r^2}\] Plug in the masses and solve the equation for r with Fg = 0

OpenStudy (anonymous):

i tried this formula but not gettin ans :(

OpenStudy (anonymous):

sure thnx...

OpenStudy (stormfire1):

Now that I think about it I'm not sure what's being asked...the last part of the question is vague. Is that the exact wording?

OpenStudy (anonymous):

yea :(

OpenStudy (stormfire1):

I think it's asking at point do the forces between the two objects cancel each other. That equation would look like this: \[\frac{m1}{r1^2} = \frac{m2}{(10m-r1)^2}\]

OpenStudy (stormfire1):

Solving for r1 in that equation will give you the point between the two objects where the forces of gravity between them are equal.

OpenStudy (anonymous):

one ques 10m what is the value of that m?

OpenStudy (stormfire1):

10 meters :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ans is not seems rit as i m getting by solving this 900/13=r1^2

OpenStudy (stormfire1):

Your final answer should be 6m btw

OpenStudy (anonymous):

u know ans is -5G/2

OpenStudy (stormfire1):

If that's their answer then we misunderstood the question...

OpenStudy (anonymous):

hey type ur ans here i have to leave byeee thnx

OpenStudy (anonymous):

there is something incomplete in ur question wat are u asking for gravitational potential or field?

OpenStudy (anonymous):

u have to find gravitational potential where the gravitational field is 0

OpenStudy (mani_jha):

the point where gravitational field is zero is x=6. Gravitational potential = -G9/6-G4/4=-G(3/2+1)=-5G/2. The second potential is -G4/4 because with respect to m=4kg, the null point is at x=4.

OpenStudy (anonymous):

 solve for x u get x=4 and other=6 later find potential by using relation -GM/r remember that field becaamre zero as o it is a VECTOR